OpenStudy (anonymous):

How do you factor out 1 + 27y^3?

OpenStudy (anonymous):

a³ + b³=( a + b )( a² – ab + b² ) 1^3+(3y)^3=(1+3y)(1^2-1*3y+(3y)^2)= (1+3y)(1-3y+9y^2))

OpenStudy (anonymous):

Thank you so much.

OpenStudy (anonymous):

You are welcome, Here is the full list [1] ( a + b )² = a² + 2ab + b² , [2] ( a – b )² = a² – 2ab + b² , [3] ( a + b ) ( a – b ) = a² – b², [4] ( a + b )³ = a³ + 3a² b + 3ab² + b³ , [5] ( a – b )³ = a ³ – 3a² b + 3ab² – b³ , [6] ( a + b )( a² – ab + b² ) = a³ + b³ , [7] ( a – b )( a ² + ab + b² ) = a³ – b³ .